- Press Release
- Dec 7, 2022
NASA Space Station On-Orbit Status 15 January 2005
SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by SpaceRef.com (copyright © 2004) to enhance access to related status reports and NASA activities.
All ISS systems continue to function nominally, except those noted previously or below. Saturday — first off-duty day for the crew (except for standard maintenance and optional task list jobs).
First thing after breakfast, FE Salizhan Sharipov terminated the bake-out cycle on filter bed 1 of the SM’s harmful impurities removal unit (BMP), moding the channel back to Purify. Later, he initiated regeneration on filter channel 2. After its completion tomorrow, the Elektron O2 generator, which uses the same vacuum vent valve, will be switched on again. [The regeneration of the air purifier filter beds is repeated every 20 days. Each bakeout to space vacuum takes about 24 hours. The BMP is currently still using the same vacuum vent valve for regeneration as the Elektron (the latter for venting hydrogen).]
CDR/SO Chiao and Sharipov performed the regular weekly 3-hr. task of thorough station cleaning. [“Uborka”, done every Saturday, includes removal of food waste products, cleaning of compartments with vacuum cleaner, wet cleaning of the Service Module (SM) dining table and other surfaces with “Fungistat” disinfectant and cleaning fan screens to avoid temperature rises.]
Sharipov transferred the accumulated data files of his BIO-5 Rasteniya-2 (“Plants-2”) plant growth experiment to the Packet laptop for subsequent downlink to the ground via Regul comm.
Because of some issues during the recent (1/10) reconfiguration of the RS OpsLAN/Ethernet SmartSwitch (BRI) from the freshly installed Wiener load, on his discretionary task list the FE was assigned the new job of downlinking previous and current BRI configuration files for review.
For his “Saturday Science” program, Leroy Chiao today performed the MFMG (Miscible Fluids in Micro Gravity) experiment’s status check of the honey and honey/water syringes, to help the Principal Investigator (PI) determine next steps for the MFMG experiments. [Crystallization of the honey was a problem last increment, so the PI wants to verify the status prior to any operations this increment. The best way to accomplish this is by photographing the honey-filled syringes for ground review. This way, investigators may decide whether or not to proceed without re-treating the honey. Background: Understanding the viscosity of molten materials is important for everything from designing laboratory experiments to industrial production of materials. One way to determine viscosity is to measure how long it takes two spheres of liquid to merge into a single spherical drop: on contact a neck will from between the two drops, increasing in thickness until the two drops become one single sphere. On Earth, gravity distorts liquid spheres, and drops are too heavy to be supported by strings. Drop distortion should not occur in the ISS’s microgravity environment, and the drops can be held on strings. To verify this technique as an accurate viscosity measurement method, the FMVM/MFMG experiment uses fluids with known viscosities: honey, corn syrup, glycerin and silicone oil.]
Leroy did the routine maintenance of the SM’s SOZh environment control & life support system, including ASU toilet facilities.
The crew performed their daily 2.5-hr. physical exercise program on TVIS treadmill, RED exerciser and VELO cycle with bungee cord load trainer. Salizhan’s daily protocol currently prescribes a four-day microcycle exercise with 1.5 hr on the TVIS (today: Day 1 of a new set) and one hour on VELO.
Leroy then transferred the daily TVIS and RED exercise data files to the MEC (medical equipment computer) for downlink, as well as the daily wristband HRM (heart rate monitor) data, followed by their erasure on the HRM storage medium.
The ISS reboost was to take place this morning at 10:18 am EST. At 7:00am, attitude control was handed over to the Russian segment motion control system (RS MCS), connected to the Progress 16 thrusters via the US-21 matching unit. Time of ignition (TIG) was at 10:18am, and handover of attitude control back to USOS at 11:30am.
The new IBM A31p ThinkPad SSC (station support computer) laptops to be installed in the SM and FGB were not deployed in the RS yesterday. The review and approval of the joint US and Russian safety data package is expected to take several more days, after which the SSC deployment schedule will be re-addressed.
Yesterday the ground performed a checkout of the new MSS (Mobile Service System) R3 software, uploaded last night to the Primary C&C MDM. Both SSRMS (Space Station Remote Manipulator System) strings and both RWS (Robotic Work Stations) checked out nominally. A problem was discovered on both the Prime and backup MBS (Mobile Base System) strings with a “watchdog” timer going into alarm. This condition would prevent the cancellation of the safing of MBS. There is no impact to the upcoming EVA #12 and the MBS is not required for mission operations until the Mission ULF1.1 timeframe. The cause of the anomaly is under investigation. The Backup and Standby C&C MDMs will continue to be loaded with the MSS R3.
The Russian Orlan EVA-12 from the Pirs Docking Compartment is scheduled for 1/26, with outside work of the ~4.5-hr. activities beginning at ~2:25am EST. A news conference to discuss details is scheduled at NASA/JSC at 1/21.
Weekly Science Update (Expedition Ten — 11th):
GASMAP: The crew was thanked for performing the GASMAP Functional Check last week. The data that the GASMAP team received on the ground assures the team that the GASMAP is in good operational shape
Human Research Facility/Workstation (HRF WS): Continuing.
Advanced Ultrasound (ADUM): Nothing new.
Hand Posture Analyzer (HPA: Nothing new.
In-Space Soldering Investigation (ISSI): Operations are complete.
Pore Formation and Mobility Investigation (PFMI): Nothing new.
Space Acceleration Measurement System (SAMS): Nothing new.
Microgravity Acceleration Measurement System (MAMS): Nothing new.
Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES): PCG-STES is performing nominally.
Protein Crystal Growth Monitoring by Digital Holographic Microscope (PromISS): Nothing new.
Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions (InSPACE): Nothing new.
Binary Colloidal Alloy Test-3 (BCAT-3): The BCAT-3 team thanked Dr. Chiao for the recent round of photographs that he took for them. Investigators are looking forward to working with the SO to photograph the surface crystals that may now be growing in BCAT Sample 10 (and possibly Samples 8 and 9.) Now that these surface crystals have been seen in Sample 10 in the remaining ground-based portion of the sample, that sample will be transferred in a cylindrical cell to a square BCAT sample cell and attempt to photograph it (as an example of what may be seen in microgravity). The color of the spots may change depending upon the angle, etc.
Renal Stone (RS): Nothing new.
Synchronized Position Hold, Engage, Reorient, Experimental Satellites (SHERES): Nothing new.
Foot/Ground Reaction Forces during Space Flight (FOOT): Nothing new.
Materials ISS Experiment (MISSE): In progress. Deployed outside on the U.S. Airlock. Nominal and collecting data.
Cellular Biotechnology Support Systems-Fluid Dynamics Investigation (CBOSS-FDI): Nothing new.
Biotechnology Specimen Temperature Controller (BSTC): Nothing new.
Yeast Group Activation Packs (Yeast GAP): Nothing new.
Earth Knowledge Acquired by Middle School Students (EarthKAM): EarthKAM operations will resume on 2/1, with the new “Firewire” cable, DCS 760 camera, and A31P laptop.
Earth Science Toward Exploration Research (ESTER): Nothing new.
Serial Network Flow Monitor (SNFM): Nothing new.
Fluid Merging Viscosity Measurement (FMVM): The FMVM team appreciated talking to Dr. Chiao and is looking forward to the experiment operations “if it becomes possible”.
Space Experiment Module (SEM): Nothing new.
Viscous Liquid Foam–Bulk Metallic Glass (Foam): Nothing new.
Effects of Prolonged Space Flight on Human Skeletal Muscle (BIOPSY): FE Sharipov was thanked for conducting the in-flight diet logging sessions over the weekend. The Biopsy team appreciates the strides the crew took to achieve our goals.
Coarsening in Solid-Liquid Mixtures-2 (CSLM-2): Planned.
Commercial Generic Bioprocessing Apparatus (CGBA): Nothing new.
Miscible Fluids in Microgravity (MFMG): Nothing new.
Educational Payload Operations (EPO): Nothing new.
Capillary Flow Experiment (CFE): Nothing new.
Crew Earth Observations (CEO): The ground has received over 6,500 CEO images to date from Increment 10. The crew was thanked for their extra efforts to expedite downlink and identification of CEO tsunami disaster imagery. The crew was also complimented on their handsome cities-at-night imagery and their efforts to correct the internal clock. An earlier image of Salalah, Oman was selected for posting on NASA’s Earth Observatory this weekend. This view illustrates the unique environment and terrain features of this ancient port city and now tourist resort.
Today’s CEO (crew earth observations) photo targets, no longer limited in the current LVLH attitude, were Tsunami damage, Nicobar Island, Thailand (looking left the island of Great Nicobar. Two minutes later look left and right for evidence of damage along the Indian Ocean coast of Thailand), Tsunami damage, Maldive Is. (pass was across the northern end of the atoll chain), Tsunami damage, Horn of Africa (this track paralleled the Somali coast just offshore. Looking left and try for overlapping images for about three minutes, for complete coverage), and Muglad Basin fans, SW Sudan (recent CEO coverage of the neighboring Salamat Basin is beginning to allow very interesting comparison of Landsat images with handheld imagery. All kinds of unexpected changes and differences of radiational response are apparent. A near-nadir mapping swath across the Muglad Basin is requested to continue this investigation of some of the planet’s least known landscapes. On the basis of handheld imagery in the last few years, it is now suspected that the modern newly understood fans are probably important for understanding underground patterns for future oil exploration in this basin).
CEO photography can be viewed and studied at the websites:
See also the website “Space Station Challenge” at:
To view the latest photos taken by the expedition 10 crew visit:
- http://spaceflight.nasa.gov/gallery/images/station/crew-10/ndxpage1.html at NASA’s Human Spaceflight website.
Expedition 10 Flight Crew Plans can be found at http://spaceflight.nasa.gov/station/timelines/
Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.
Upcoming Key Events:
- ISS reboost (delta-V: 4 m/s) — 1/15/05, at 10:18 am EST;
- EVA-12 — 1/26/05 (Eastern)
- Progress 16P undocking & destructive reentry — 2/27/05;
- Progress 17P launch — 2/28/05.
- EVA-13 — 3/25/05;
- Soyuz 10 S launch — 4/15/05;
- Soyuz 9S undock — 4/25/05 (after 193 days on orbit, 191 days on board ISS).
ISS Altitude History
Apogee height — Mean Altitude — Perigee height
For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html. In addition, information on International Space Station sighting opportunities can be found at http://spaceflight.nasa.gov/realdata/sightings/ on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at http://science.nasa.gov/temp/StationLoc.html at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at http://www.spaceref.com/iss/tracking.html.